The perimeter of a triangle is 38 feet. the length is 17 feet less than five times the width. What are the dimensions of the rectangle?
Answer by Alan3354(69443)

The perimeter of a triangle is 38 feet. the length is 17 feet less than five times the width. What are the dimensions of the rectangle?
------------------
I went past this one a few times because it says "triangle", then it says "rectangle." I think it's a rectangle.
---------------
P = 2L + 2W = 38 (perimeter = 2*width + 2*length)
L + W = 19
L = 5W - 17
Sub for L
5W-17 + W = 19
6W = 36
W = 6 feet
L = 13 feet